The best laboratory vacuum systems can pump down to as few as 1x10^9 molecules per cubic meter of gas. Calculate the corresponding pressure, in atmospheres, assuming a temperature of 298K
1
Expert's answer
2016-12-10T03:12:10-0500
P = nKT = 1*109*1.38*10-23*298 =4.11*10-11 (Pa) = 4*10-16 (atm)